# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ 3.2 Working on the sphere
Astronomical observations can be viewed as images of a three-dimensional space projected onto the inside of a two-dimensional surface of a sphere at a given distance with the observer at its centre. Therefore, galaxy position measurements, most observational biases and some systematics are naturally expressed in spherical coordinates on the sky. For this reason, the forward model presented in this work is expressed in the same geometry.
The KiDS-SBI forward simulations use GLASS (Tessore et al., 2023) to model the underlying large-scale structure. Within GLASS, we choose to model the large-scale structure through log-normal random matter fields within non-overlapping concentric shells centred at the observer. Therefore, a given shell is a comoving volume which spans the full sky and has a finite width along the line-of-sight from one redshift, $z_{i}$, to another, $z_{i+1}$, such that $i\in\{1,2,3,...,N_{\mathrm{shells}}\}$. We define the matter weight function $W^{(i)}(z)$ describing the distribution along the line-of-sight of the comoving volume spanned by a given shell as (Tessore et al., 2023),
@@
A given shell should not be thicker than a few hundred $\mathrm{Mpc}$ of comoving distance along the line-of-sight for log-normal random fields to be an accurate enough representation of the distribution of large-scale structure (Xavier et al., 2016; Hall & Taylor, 2022; Tessore et al., 2023). Otherwise, discretisation effects would smooth away large-scale structure (Tessore et al., 2023). Additionally, one needs to be mindful of computational resources, because the number required of angular power spectra between the matter fields within each shell scales as $2N_{\mathrm{shells}}$. Simultaneously, the simulations’ run-time for a given set of cosmological parameters increases with $N_{\mathrm{shells}}$. To balance these factors, we aim to have as many shells along the line-of-sight as computationally feasible.
@@ 3.3 Non-Limber projection
Although the three-dimensional matter power spectrum characterises all cosmological dependence of the matter fields in our simulation, for it to define the cosmology dependence in the simulated matter fields, it has to be projected according to the geometry of the shells considered in Sect. 3.2. To determine the correlations between the matter fields of comoving spherical shells, we define the two-dimensional angular matter power spectra for a given cosmology, $\mathbf{\Theta}$, and a given shell combination $(ij)$ as follows
@@
where $\,{}_{0}\!\!\operatorname{Y}_{\ell m}(\bm{\theta})$ are the spherical harmonics for a spin value of 0 as a function of the two-dimensional spatial sky position $\bm{\theta}$. The projected matter field, $\delta^{(i)}(\bm{\theta};\mathbf{\Theta})$, can in turn be defined with respect to the underlying three-dimensional matter overdensity field, $\delta(\bm{\theta},z;\mathbf{\Theta})$, as
@@
The three-dimensional matter overdensity field’s cosmology dependence can be quantified as follows
@@
where $\delta(\mathbf{k},z;\mathbf{\Theta})$ is the Fourier transform of the two-dimensional matter overdensity field at $z$, $\delta(\bm{\theta},z,;\mathbf{\Theta})$, while $\delta_{\mathrm{D}}(\mathbf{k}-\mathbf{k}^{\prime})$ is the Dirac delta function.
If we combine the relations in Eqs. (4) to (7) and, for the sake of simplicity, assume that $\Omega_{k}=0$, i.e. $f_{k}(z;\mathbf{\Theta})=\chi(z;\mathbf{\Theta})$, we can characterise the cosmology dependence of $C_{\delta\delta}^{(ij)}(\ell;\mathbf{\Theta})$ directly with respect to the three-dimensional matter power spectrum, $P_{\delta,\mathrm{nl}}(k,z,z^{\prime};\mathbf{\Theta})$, using the following relation
@@
where $\chi^{\prime}\equiv z[\chi^{\prime}]$, $j_{\ell}(k\chi)$ are spherical Bessel functions of order $\ell$, and $\ell\in\mathbb{Z}^{0+}$. We can simplify Eq. (8) by making the geometric approximation, i.e. taking the geometric mean of two equal-time power spectra (Castro et al., 2005; Kitching & Heavens, 2017; Kilbinger et al., 2017; de la Bella et al., 2021), such that
@@
Figure 4: The two-dimensional angular matter power spectra, $C_{\delta\delta}^{(ij)}(\ell)$, projected with Levin describing the correlations between the large-scale structure within a set of 19 concentric shells spaced along the line-of-sight as shown in Fig. 2. The left panel shows the autocorrelations of all shells, i.e. $|i-j|=0$. The right panel shows the correlations of each shell with its nearest neighbour, i.e. $|i-j|=1$ ($|i-j|>1$ are not shown as they do not have a large effect on matter fields within each shell, see Tessore et al. 2023). The colour of each line is given by the mean redshift of the $i^{\mathrm{th}}$ bin, $\overline{z}^{(i)}$. The underlying linear three-dimensional matter power spectrum is based on flat $\Lambda$CDM assuming $\Omega_{\mathrm{c}}=0.05$, $\Omega_{\mathrm{b}}=0.28$, $\sigma_{8}=0.79$, $S_{8}=0.84$ and $H_{0}=67$ km s-1 Mpc-1 calculated using CAMB (Lewis et al., 2000; Lewis & Challinor, 2002; Howlett et al., 2012), while the non-linear contribution is calculated using HMcode-2016 (Mead et al., 2015, 2016) assuming $A_{\mathrm{bary}}=3.1$. The non-Limber projection done by Levin assumes $\ell_{\mathrm{max,nL}}=1200$ which causes the $C_{\delta\delta}^{(ij)}$ with $|i-j|>0$ go to a value of zero for $\ell>\ell_{\mathrm{max,nL}}$.
Additionally, it is common to make the so-called Limber approximation (Limber, 1953; Kaiser, 1992) which consists in approximating a spherical Bessel function as a single peak or delta function at $k\chi=\ell+1/2$. This approximation is reasonably accurate in the auto-correlations for broad window functions and high angular modes, $C_{\delta\delta}^{(ii)}(\ell;\mathbf{\Theta})$, but depending on the thickness of the shells, the Limber approximation can lead to substantial biases, especially, at large scales. However, when making the Limber approximation, all off-diagonal angular power spectra, $C_{\delta\delta}^{(ij)}$, where $i\neq j$, will be zero as long as their associated weight functions, $W^{(i)}(z[\chi];\mathbf{\Theta})$, are non-overlapping, as is the case for the weights constructed in Eq. (2). This can be seen by looking at the full Taylor expansion of Eq. (8) as given by (LoVerde & Afshordi, 2008)
@@
where each apostrophe, , denotes a partial derivative with respect to $\chi$, i.e. $\partial/\partial\chi$, while we also omit the explicit cosmological dependence of the power spectrum and the weight functions here for the sake of clarity. The non-Limber integral therefore also contributes to the autocorrelations, where $|i-j|=0$, particularly for small $\ell$ and the narrow weight functions considered here. If not considered this can induce a large bias in the angular power spectra of the shells which could significantly bias cosmological constraints from KiDS-1000. At the same time, we find that for the cross-correlations of adjacent shells, such that $|i-j|=1$, the angular power can be of the order of $10^{-1}C_{\delta\delta}^{(ii)}$ or less. In this case, the simulated $\delta^{(i)}(\bm{\theta};\mathrm{\Theta})$ would have a percent-level bias while being void of any large-scale structure along the line-of-sight beyond the scales of individual shells. To avoid this, we choose to calculate the full non-Limber projection as given by Eq. (8). The correlations between shell pairs which are not immediate neighbours, i.e. $|i-j|>1$, contribute less than $0.1\%$ to the overall angular power of a shell, so their effect is undetectable for KiDS.
@@
The module allows one to divide the domain of the angular power spectra into three regions: a non-Limber domain, a second-order extended Limber domain and a Limber domain. In this work, we choose to split the calculation of $C_{\delta\delta}^{(ij)}(\ell)$ into two domains: non-Limber ($1<\ell\leq\ell_{\mathrm{max,nL}}$) and second-order extended Limber given in Eq. (10) for $\ell_{\mathrm{max,nL}}<\ell\leq 30,000$. Over the former domain, Levin numerically integrates the expression shown in Eq. (8) for the given matter power spectrum and the weights. Over the latter domain, we perform the numerical integration of Eq. (10) up to the second-order term in $(\ell+1/2)^{-1}$. This reduces the computational resources needed for the integration without any substantial loss of accuracy as long as $\ell_{\mathrm{max,nL}}$ is sufficiently large, since the residuals scale with $(\ell+1/2)^{-4}$ (LoVerde & Afshordi, 2008). With the shells shown in Fig. 2 weighted with Eq. (2), we obtain angular power spectra, $C_{\delta\delta}^{(ij)}(\ell;\mathbf{\Theta})$, shown in Fig. 4.
@@ 3.4 Log-normal matter field simulations
With the input angular power spectra for a given cosmology computed as described in Appendix A, the next step in the forward simulations involves sampling random matter fields consistent with these correlations. These fields are generated with GLASS (Tessore et al. 2023). This framework allows one to create efficient and accurate random field simulations of large-scale structure matter fields as well as the associated weak gravitational lensing signals. Within the GLASS suite, many different choices can be made, while it also allows one to sample the angular positions, redshifts, shapes and shears of galaxies. Nevertheless, in this analysis, we only choose to use it to simulate the underlying matter overdensity fields, $\delta^{(i)}(\bm{\theta};\mathbf{\Theta})$, as well as the associated convergence fields for each shell, $\kappa^{(i)}(\bm{\theta};\mathbf{\Theta})$. These are then integrated along the line-of-sight to give the convergence field for a source within a given shell, $\kappa^{(i)}$.
@@
Log-normal fields within GLASS are based on Gaussian random fields. One can obtain these by taking the angular power spectra of the shells and assuming that they describe the correlations between log-normal fields, $\delta_{\mathrm{log}}^{(i)}(\bm{\theta};\mathbf{\Theta})$, as follows
@@
one can then sample $\delta_{\mathrm{G}}^{(i)}(\bm{\theta};\mathbf{\Theta})$ for each shell. To achieve this with numerical methods, the two-dimensional spherical shell of each layer is discretised using HEALPix pixels (Górski et al., 2005). This means that $\bm{\theta}$ is discretised such that it can be mapped to a linearised variable given by $\theta_{m}$ where $m\in\{1,2,3,...,12N_{\mathrm{side}}^{2}\}$ and $N_{\mathrm{side}}$ is the HEALPix resolution parameter. To obtain the final matter overdensity fields, we use the following expression (Coles & Jones, 1991; Kayo et al., 2001; Hilbert et al., 2011; Xavier et al., 2016),
@@
The layers of concentric log-normal matter overdensity fields within each shell give a full description of the large-scale structure. The resolution of this structure will of course be limited by the size of the shells along the line-of-sight, while being limited by the resolution of the discretised two-dimensional pixels on the sphere. However, as discussed with regards to the resolution of shells in Sect. 3.2, if one chooses sufficiently large $\ell_{\mathrm{max}}$ up to which the input angular power spectra are calculated, one can accurately sample $\delta_{\mathrm{log}}^{(i)}(\theta_{m};\mathbf{\Theta})$ up to an $N_{\mathrm{side}}\sim 0.5\ell_{\mathrm{max}}$ (Leistedt et al., 2013; Alonso et al., 2019).
As we intend to make use of these simulations for weak gravitational lensing, the next step in the forward simulations is to model how the image of galactic sources within a given shell will be distorted by weak gravitational lensing due to the matter overdensities the light encounters along the line-of-sight to the observer. Here, the geometry of the simulations aids us again, since the concentric volumes in comoving distance that each shell makes up intrinsically allow one to trace all possible light-cones emanating from the observer.
We start from the definition of the convergence for a source located at $\bm{\theta}$ and $z$ along a continuous line-of-sight under the Born approximation (Bartelmann & Schneider, 2001),
@@
where $\Omega_{\mathrm{m}}$ is the matter density fraction at $z=0$, and $\Omega_{\mathrm{m}}\in\mathbf{\Theta}$. We then discretise the continuous two-dimensional matter overdensity field along over $\bm{\theta}$ using HEALPix pixels, while also discretising along the line-of-sight using Eq. (6) and defining a mean redshift, $\overline{z}^{(i)}$, for a given weighted shell as follows
$\overline{\mathit{z}}^{(i)}(\mathbf{\Theta})=\frac{\int\mathrm{d}\mathit{z}\,\mathit{z}\,\mathit{W}^{(i)}(\mathit{z};\mathbf{\Theta})}{\int\mathrm{d}\mathit{z}\,\mathit{W}^{(i)}(\mathit{z});\mathbf{\Theta}}\,.$ (15)
To accelerate the calculation of the discretised convergence fields for each shell, $\kappa^{(i)}$, within GLASS, it makes use of the fact that for all Robertson-Walker space-times, transverse comoving distances will scale in such a way that one can write down a recurrence relation from one interval to the next along the line-of-sight (Schneider, 2016). This gives a recurrence relation for the convergence field, $\kappa^{(i)}$, within shells where $i\geq 2$, which only depends on $\kappa^{(i-1)}$, $\kappa^{(i-2)}$ and $\delta^{(i-1)}$ (see Tessore et al. 2023 for details). Such a relation allows one to calculate the convergence fields for all shells, while only holding three fields in memory at a given time. This greatly reduces the amount of computational resources needed per realisation, so the large amount of simulations at different cosmologies, $\mathbf{\Theta}$, needed for this analysis can be obtained with fewer resources when compared to forward models in previous analysis (see Sect. 5.1).
We now have all the ingredients to sample galaxies and to shear their shapes, while simultaneously having the flexibility to add any field-level systematics which may be relevant.

### v2

@@ 3.2 Working on the sphere
Astronomical observations can be viewed as images of a 3D space projected onto the inside of a 2D surface of a sphere at a given distance with the observer at its centre. Therefore, galaxy position measurements, most observational biases and some systematics are naturally expressed in spherical coordinates on the sky. For this reason, the forward model presented in this work is expressed in the same geometry.
The KiDS-SBI forward simulations use GLASS (Tessore et al., 2023) to model the underlying large-scale structure. Within GLASS, we chose to model the large-scale structure through log-normal random matter fields within non-overlapping concentric shells centred at the observer. Therefore, a given shell is a comoving volume which spans the full sky and has a finite width along the line of sight from one redshift, $z_{i}$, to another, $z_{i+1}$, such that $i\in\{1,2,3,...,N_{\mathrm{shells}}\}$. We define the matter weight function $W^{(i)}(z)$ describing the distribution along the line of sight of the comoving volume spanned by a given shell as (Tessore et al., 2023),
@@
A given shell should not be thicker than a few hundred $\mathrm{Mpc}$ of comoving distance along the line of sight for log-normal random fields to be an accurate enough representation of the distribution of large-scale structure (Xavier et al., 2016; Hall & Taylor, 2022; Tessore et al., 2023). Otherwise, discretisation effects would smooth away large-scale structure (Tessore et al., 2023). Additionally, we need to be mindful of computational resources, because the number required of angular power spectra between the matter fields within each shell scales as $2N_{\mathrm{shells}}$. Simultaneously, the simulation run-time for a given set of cosmological parameters increases with $N_{\mathrm{shells}}$. To balance these factors, we aim to have as many shells along the line of sight as computationally feasible.
@@ 3.3 Non-Limber projection
Although the 3D matter power spectrum characterises all cosmological dependence of the matter fields in our simulation, for it to define the cosmology dependence in the simulated matter fields, it has to be projected according to the geometry of the shells considered in Sect. 3.2. To determine the correlations between the matter fields of comoving spherical shells, we define the 2D angular matter power spectra for a given cosmology, $\mathbf{\Theta}$, and a given shell combination $(ij)$ as follows
@@
where $\,{}_{0}\!\!\operatorname{Y}_{\ell m}(\bm{\theta})$ are the spherical harmonics for a spin value of 0 as a function of the 2D spatial sky position $\bm{\theta}$. The projected matter field, $\delta^{(i)}(\bm{\theta};\mathbf{\Theta})$, can in turn be defined with respect to the underlying 3D matter overdensity field, $\delta(\bm{\theta},z;\mathbf{\Theta})$, as
@@
The 3D matter overdensity field’s cosmology dependence can be quantified as follows
@@
where $\delta(\mathbf{k},z;\mathbf{\Theta})$ is the Fourier transform of the 3D matter overdensity field at $z$, $\delta(\bm{\theta},z,;\mathbf{\Theta})$, while $\delta_{\mathrm{D}}(\mathbf{k}-\mathbf{k}^{\prime})$ is the Dirac delta function.
If we combine the relations in Eqs. (4) to (7) and, for the sake of simplicity, assume that $\Omega_{k}=0$, specifically $f_{k}(z;\mathbf{\Theta})=\chi(z;\mathbf{\Theta})$, we can characterise the cosmology dependence of $C_{\delta\delta}^{(ij)}(\ell;\mathbf{\Theta})$ directly with respect to the 3D matter power spectrum, $P_{\delta,\mathrm{nl}}(k,z,z^{\prime};\mathbf{\Theta})$, using the following relation
@@
where $\chi^{\prime}\equiv z[\chi^{\prime}]$, $j_{\ell}(k\chi)$ are spherical Bessel functions of order $\ell$, and $\ell\in\mathbb{Z}^{0+}$. We can simplify Eq. (8) by making the geometric approximation, that is taking the geometric mean of two equal-time power spectra (Castro et al., 2005; Kitching & Heavens, 2017; Kilbinger et al., 2017; de la Bella et al., 2021), such that
@@
Figure 4: 2D angular matter power spectra, $C_{\delta\delta}^{(ij)}(\ell)$, projected with Levin describing the correlations between the large-scale structure within a set of 19 concentric shells spaced along the line of sight as shown in Fig. 2. The top panel shows the autocorrelations of all shells, i.e. $|i-j|=0$. The bottom panel shows the correlations of each shell with its nearest neighbour, i.e. $|i-j|=1$ ($|i-j|>1$ are not shown as they do not have a large effect on matter fields within each shell, see Tessore et al. 2023). The colour of each line is given by the mean redshift of the $i^{\mathrm{th}}$ bin, $\overline{z}^{(i)}$. The underlying linear 3D matter power spectrum is based on flat $\Lambda$CDM assuming $\Omega_{\mathrm{c}}=0.05$, $\Omega_{\mathrm{b}}=0.28$, $\sigma_{8}=0.79$, $S_{8}=0.84$ and $H_{0}=67$ km s-1 Mpc-1 calculated using CAMB (Lewis et al., 2000; Lewis & Challinor, 2002; Howlett et al., 2012), while the non-linear contribution is calculated using HMcode-2016 (Mead et al., 2015, 2016) assuming $A_{\mathrm{bary}}=3.1$. The non-Limber projection done by Levin assumes $\ell_{\mathrm{max,nL}}=1200$ which causes the $C_{\delta\delta}^{(ij)}$ with $|i-j|>0$ go to a value of zero for $\ell>\ell_{\mathrm{max,nL}}$.
Additionally, it is common to make the so-called Limber approximation (Limber, 1953; Kaiser, 1992) which consists in approximating a spherical Bessel function as a single peak or delta function at $k\chi=\ell+1/2$. This approximation is reasonably accurate in the auto-correlations for broad window functions and high angular modes, $C_{\delta\delta}^{(ii)}(\ell;\mathbf{\Theta})$, but depending on the thickness of the shells, the Limber approximation can lead to substantial biases, especially, at large scales (Lemos et al., 2017). However, when making the Limber approximation, all off-diagonal angular power spectra, $C_{\delta\delta}^{(ij)}$, where $i\neq j$, will be zero as long as their associated weight functions, $W^{(i)}(z[\chi];\mathbf{\Theta})$, are non-overlapping, as is the case for the weights constructed in Eq. (2). This can be seen by looking at the full Taylor expansion of Eq. (8) as given by (LoVerde & Afshordi, 2008)
@@
where each apostrophe, , denotes a partial derivative with respect to $\chi$, specifically $\partial/\partial\chi$, while we also omit the explicit cosmological dependence of the power spectrum and the weight functions here for the sake of clarity. The non-Limber integral therefore also contributes to the autocorrelations, where $|i-j|=0$, particularly for small $\ell$ and the narrow weight functions considered here. If not considered this can induce a large bias in the angular power spectra of the shells which could significantly bias cosmological constraints from KiDS-1000. At the same time, we find that for the cross-correlations of adjacent shells, such that $|i-j|=1$, the angular power can be of the order of $10^{-1}C_{\delta\delta}^{(ii)}$ or less. In this case, the simulated $\delta^{(i)}(\bm{\theta};\mathrm{\Theta})$ would have a percent-level bias while being void of any large-scale structure along the line of sight beyond the scales of individual shells. To avoid this, we chose to calculate the full non-Limber projection as given by Eq. (8). The correlations between shell pairs which are not immediate neighbours, that is $|i-j|>1$, contribute less than $0.1\%$ to the overall angular power of a shell, so their effect is undetectable for KiDS.
@@
The module allows us to divide the domain of the angular power spectra into three regions: a non-Limber domain, a second-order extended Limber domain and a Limber domain. In this work, we chose to split the calculation of $C_{\delta\delta}^{(ij)}(\ell)$ into two domains: non-Limber ($1<\ell\leq\ell_{\mathrm{max,nL}}$) and second-order extended Limber given in Eq. (10) for $\ell_{\mathrm{max,nL}}<\ell\leq 30,000$. Over the former domain, Levin numerically integrates the expression shown in Eq. (8) for the given matter power spectrum and the weights. Over the latter domain, we perform the numerical integration of Eq. (10) up to the second-order term in $(\ell+1/2)^{-1}$. This reduces the computational resources needed for the integration without any substantial loss of accuracy as long as $\ell_{\mathrm{max,nL}}$ is sufficiently large, since the residuals scale with $(\ell+1/2)^{-4}$ (LoVerde & Afshordi, 2008). With the shells shown in Fig. 2 weighted with Eq. (2), we obtain angular power spectra, $C_{\delta\delta}^{(ij)}(\ell;\mathbf{\Theta})$, shown in Fig. 4. We note that, in Fig. 4, the Limber extension to second order from Eq. (10) calculated for $\ell>1200$ reduces roughly to the Limber approximation as the higher-order terms have a vanishing contribution to the overall angular power spectrum. In fact, in the cross-correlations between shells where $|i-j|>0$, the second order term in Eq. (10) becomes vanishingly small for these values of $\ell$. Consequently, we effectively apply a scale cut in the matter shell cross-correlations for $\ell>1200$. We find that this does not bias our analysis, as the relative contribution cross-correlations to the total $C^{(ij)}_{\delta\delta}(\ell)$ of a given shell is less than $0.1\%$ for $\ell\geq 1200$ for all redshifts. This ensures sub-percent level accuracy of the input matter fields across all scales considered in this analysis.
@@ 3.4 Log-normal matter field simulations
With the input angular power spectra for a given cosmology computed as described in Appendix A, the next step in the forward simulations involves sampling random matter fields consistent with these correlations. These fields are generated with GLASS (Tessore et al. 2023). This framework allows us to create efficient and accurate random field simulations of large-scale structure matter fields as well as the associated weak gravitational lensing signals. Within the GLASS suite, many different choices can be made, while it also allows us to sample the angular positions, redshifts, shapes and shears of galaxies. Nevertheless, in this analysis, we only chose to use it to simulate the underlying matter overdensity fields, $\delta^{(i)}(\bm{\theta};\mathbf{\Theta})$, as well as the associated convergence fields for each shell, $\kappa^{(i)}(\bm{\theta};\mathbf{\Theta})$. These are then integrated along the line of sight to give the convergence field for a source within a given shell, $\kappa^{(i)}$.
@@
Log-normal fields within GLASS are based on Gaussian random fields. We can obtain these by taking the angular power spectra of the shells and assuming that they describe the correlations between log-normal fields, $\delta_{\mathrm{log}}^{(i)}(\bm{\theta};\mathbf{\Theta})$, as follows
@@
we can then sample $\delta_{\mathrm{G}}^{(i)}(\bm{\theta};\mathbf{\Theta})$ for each shell. To achieve this with numerical methods, the 2D spherical shell of each layer is discretised using HEALPix pixels (Górski et al., 2005). This means that $\bm{\theta}$ is discretised such that it can be mapped to a linearised variable given by $\theta_{m}$ where $m\in\{1,2,3,...,12N_{\mathrm{side}}^{2}\}$ and $N_{\mathrm{side}}$ is the HEALPix resolution parameter. To obtain the final matter overdensity fields, we use the following expression (Coles & Jones, 1991; Kayo et al., 2001; Hilbert et al., 2011; Xavier et al., 2016),
@@
The layers of concentric log-normal matter overdensity fields within each shell give a full description of the large-scale structure. The resolution of this structure will of course be limited by the size of the shells along the line of sight, while being limited by the resolution of the discretised 2D pixels on the sphere. However, as discussed with regards to the resolution of shells in Sect. 3.2, if we chose a sufficiently large $\ell_{\mathrm{max}}$ value up to which the input angular power spectra are calculated, we can accurately sample $\delta_{\mathrm{log}}^{(i)}(\theta_{m};\mathbf{\Theta})$ up to an $N_{\mathrm{side}}\sim 0.5\ell_{\mathrm{max}}$ (Leistedt et al., 2013; Alonso et al., 2019).
As these simulations were performed prior to carrying out the weak gravitational lensing study, the next step in the forward simulations is to model how the image of galactic sources within a given shell will be distorted by weak gravitational lensing due to the matter overdensities the light encounters along the line of sight to the observer. Here, the geometry of the simulations aids us again, since the concentric volumes in comoving distance that each shell makes up intrinsically allow us to trace all possible light cones emanating from the observer.
We start from the definition of the convergence for a source located at $\bm{\theta}$ and $z$ along a continuous line of sight under the Born approximation (Bartelmann & Schneider, 2001),
@@
where $\Omega_{\mathrm{m}}$ is the matter density fraction at $z=0$, and $\Omega_{\mathrm{m}}\in\mathbf{\Theta}$. We then discretise the continuous 2D matter overdensity field along over $\bm{\theta}$ using HEALPix pixels, while also discretising along the line of sight using Eq. (6) and defining a mean redshift, $\overline{z}^{(i)}$, for a given weighted shell as follows
$\overline{\mathit{z}}^{(i)}(\mathbf{\Theta})=\frac{\int\mathrm{d}\mathit{z}\,\mathit{z}\,\mathit{W}^{(i)}(\mathit{z};\mathbf{\Theta})}{\int\mathrm{d}\mathit{z}\,\mathit{W}^{(i)}(\mathit{z};\mathbf{\Theta})}\,.$ (15)
To accelerate the calculation of the discretised convergence fields for each shell, $\kappa^{(i)}$, within GLASS, it makes use of the fact that for all Robertson-Walker space-times, transverse comoving distances will scale in such a way that we can write down a recurrence relation from one interval to the next along the line of sight (Schneider, 2016). This gives a recurrence relation for the convergence field, $\kappa^{(i)}$, within shells where $i\geq 2$, which only depends on $\kappa^{(i-1)}$, $\kappa^{(i-2)}$ and $\delta^{(i-1)}$ (see Tessore et al. 2023 for details). Such a relation allows us to calculate the convergence fields for all shells, while only holding three fields in memory at a given time. This greatly reduces the amount of computational resources needed per realisation, so the large amount of simulations at different cosmologies, $\mathbf{\Theta}$, needed for this analysis can be obtained with fewer resources when compared to forward models in previous analysis (see Sect. 5.1). Once we have all the ingredients to sample galaxies and to shear their shapes, while simultaneously having the flexibility to add any relevant field-level systematics, we can proceed to the next phase of the study.